Ron's building, which was used in his business, was destroyed in a fire. Ron's adjusted basis in the building was $210,000, and its FMV was $330,000. Ron filed an insurance claim and was reimbursed $300,000. In that same year, Ron invested $240,000 of the insurance proceeds in another business building. Ron will recognize gain of

Group of answer choices

$30,000

$0

$60,000

$90,000
Answer #1

Answer)

a)Insurance proceeds=$300,000
Adjusted basis of old building=$210,000
1)Gain realized= Insurance proceed- Adjusted basis of old building
=300,000- 210,000=$90,000

2)Gain Recognized= Insurance proceeds-Insurance proceeds Reinvested
=300,000 - 240,000
=60,000

Gain Recognized=minimum of gain realized and gain recognized
=min(90000, 60,000= 60,000

Answer is $60,000